Explain the fact: A fixed force is applied for a certain interval of time on 2 different bodies of different masses,initially at rest,the lighter body acquires greater speed than the heavier body. At the end of time observations show that each body acquires the same momentum.Thus the same force for the same time causes the same change in momentum for different bodies.


Reference: NCERT PHYSICS TEXTBOOK-HIGHER SECONDARY FIRST YEAR-CHAPTER:LAWS OF MOTION

Grade:12

1 Answers

Puneet Mittal AskiitiansExpert-IITD
13 years ago

Hi Akshaya,

When u apply same force F on 2 bodies of Masses m1 and m2 then both will have different accelarations a1 and a2.

Let m1 less than m2 , then a1 will be greater than a2 ( because F = m1a1 and F = m2a2 )

Thus ligther body acquires greater speed than the heavier body.

At end of time t , v1 = 0 + 1/2 * a1 * t2

 v1 = 1/2*(F/m1)*t2  or m1v1 = 1/2*F*t2

Similarly at end of time t , v2 = 1/2*a2*t2

v2 = 1/2*(F/m2)*t2  or m2v2 = 1/2*F*t2

Hence u can see that both acquires the same momentum at the end because they have same accelaration or same force acting upon them ( both bodies will have diff velocities at the end as they have diff masses )

Because force is change in momentum of a body so same force will always cause same change in momentum.

Please feel free to ask as many questions you have.

Puneet